what is x in ; 1/3x + 1/2 = 11/12
Found 2 solutions by unlockmath, Aaniya:
Answer by unlockmath(1688)

Hello,
First, I'm assuming this is written as:
1/3(x) + 1/2 = 11/12
If so then multiply by 12 to get:
4x+6=11
Subtract 6 and divide by 4 to get:
x=5/4
Make sense?
RJ
www.math-unlock.com

Answer by Aaniya(63)